OpenStudy (anonymous):

If f(x) = 3x2 - x, find f(-2). 10 14 38

OpenStudy (anonymous):

we have to put (-2) in place of x n given equation, f(-2)=3*(-2)^2-(-2) =3*4+2 =12+2 =14

OpenStudy (anonymous):

ahh thank you :)